# Conformally rigid graphs

### v1

@@ Abstract.
Conformal rigidity requires an extraordinary amount of symmetry in $G$. Every edge-transitive graph is conformally rigid. We prove that every distance-regular graph, and hence every strongly-regular graph, is conformally rigid. Certain special graph embeddings can be used to characterize conformal rigidity. Cayley graphs can be conformally rigid but need not be, we prove a sufficient criterion. We also find a small set of conformally rigid graphs that do not belong into any of the above categories; these include the Hoffman graph, the crossing number graph 6B and others. Conformal rigidity can be certified via semidefinite programming, we provide explicit examples.
@@ 1.2. Motivation
A final motivation comes from a recently proposed framework for Graph Sparsification due to Babecki and the authors [1]. In this framework, the hypercube graph cannot be further sparsified. We were wondering how general this result is which is what lead us to the notion of conformally rigid graphs. If (up to scaling) $\mathbbm{1}$ is the unique weight that maximizes $\lambda_{2}(w)$, then $G$ cannot be sparsified in any setting of [1]. This led us to wonder if all Cayley graphs have this property. As we will see in this paper, all Cayley graphs do not have this property. This line of investigation led us to the notion of conformal rigidity. It is an interesting question to understand which conformally rigid graphs have unique maximizers for one or both of $\lambda_{2}(w)$ and $\lambda_{n}(w)$.
@@ 1.3.1. The optimization perspective
The spectral width minimization problem looks similar to conformal rigidity and so we pause to point out an important difference. In our setup we are solving
@@ 2.3. Distance-regular graphs
Note that, in particular, every strongly-regular graph is distance-regular. Strongly-regular graphs are distance-regular graphs of diameter two. For example, the Shrikhande graph and its complement (see Fig. 1) are strongly-regular. The proof of Theorem 2.2 uses structural theory of distance-regular graphs (see [11]) and the duality theory of semidefinite programs which are developed in § 3 and § 4.
@@ 2.4. Cayley Graphs
A natural class of graphs with a great deal of symmetry are Cayley graphs $\Gamma(G,S)$: these arise when combining a finite group $G$ with a generating set $S\subset G$. The Cayley graph $\Gamma(G,S)$ has vertex set $V=G$ and the edge set $E=\left\{(g,s\circ g):g\in G,s\in S\right\}.$ Since we only deal with undirected graphs, we want the edge set to be symmetric which is achieved by requiring a symmetry in the generating set $S$, meaning $s\in S\implies s^{-1}\in S$. We always assume this.
Cayley graphs need not be conformally rigid. The smallest example we found is shown on the left in Fig. 5. This unweighted graph has $\lambda_{2}=2$ and $\lambda_{6}=5$. As it turns out, neither of these eigenvalues are conformally optimal: by increasing the weights on the 3 edges connecting the two triangles (while decreasing the weight on the other edges to maintain the normalization), we can increase $\lambda_{2}$ and attain $\lambda_{2}(w)=18/7>2$. Likewise, by setting the three edge weights connecting the two triangles to 0 (and increasing the remaining weights uniformly) we can decrease the largest eigenvalue to $\lambda_{6}(w)=4.5<5$ (and thus, in particular, can achieve values smaller than 5 for small positive weights by continuity of the eigenvalues). On the other hand, the circulant graph shown on the right of Fig. 5 is conformally rigid but not edge-transitive. Its conformal rigidity is implied by the following theorem.
@@
Figure 5. Left: a Cayley graph that is not conformally rigid. Right: a Cayley graph on $\mathbb{Z}_{18}$ (generated by $S=\left\{-5,-1,1,5\right\}$) that is conformally rigid and not edge-transitive.
@@
For the circulant graph on the right in Fig. 5, $V={\mathbb{Z}}_{18}=\{0,1,2,\ldots,17\}$ and $S=\{1,-1,5,-5\}$. We choose
@@ 2.5. Circulants
Denote a circulant with $n$ vertices and generators $S$ by $C_{n}(S)$. It is typical to list only one of $s$ or $-s$ in $S$, with the assumption that the other one is included. The set of vertices of $C_{n}(S)$ is $V={\mathbb{Z}}_{n}=\left\{0,1,2,\dots,n-1\right\}$ and its edges are $(v,v+s)$, as we vary over all $v\in V$ and $s\in S$ (cyclically interpreted). Since circulants are Cayley graphs, Theorem 2.3 applies to them. We start with some examples.
@@
Circulants are rich and rather nontrivial. In particular, understanding when a circulant is edge-transitive appears to be a nontrivial problem, no definite characterization appears to be known [2, 5, 6, 25, 28, 29]. The first goal of this section is to prove that there are infinitely many circulants that are not conformally rigid.
@@
    Somewhat unrelated, but of independent interest in this context: is it possible to characterize when $C_{n}(\left\{a,b\right\})$ is edge-transitive in terms of $n,a,b$?
@@ 2.6. Sporadic outliers
The graphs in Theorem 2.6 were all proven to be conformally rigid by identification of a suitable dual semidefinite programming certificate (see § 6). However, we also found one other example, the non-Cayley vertex-transitive graph (20,10) on $n=20$ vertices, that appeared to be not edge-transitive, not distance-regular and very likely conformally rigid (beyond any reasonable numerical doubt); however, we were unable to find a certificate that would rigorously prove its conformal rigidity.
@@ 3. Semidefinite programming certificates for conformal rigidity
Assume that $G=([n],E)$ is an undirected connected graph with positive edge weights $w_{ij}$. Let $\mathcal{S}_{+}^{n}$ be the cone of $n\times n$ real symmetric psd matrices. We use $X\succeq 0$ to denote that $X\in\mathcal{S}^{n}_{+}$. For $A,B\in\mathcal{S}^{n}_{+}$, $A\succeq B$ implies that $A-B\succeq 0$.
@@
Problem (8) has an optimal solution since its objective function is continuous and its feasible region is compact. Therefore, (9) also has an optimal solution. If $w^{*}$ is optimal for (9), then $\lambda_{2}(w^{*})=1$ and hence there is no harm in optimizing over $\lambda_{2}(w)\geq 1$. The two problems are off by a scaling; $w^{*}$ is optimal for (9) if and only if $\lambda_{2}(\mathbbm{1})w^{*}$ is optimal for (8), and $\lambda_{2}^{*}=\lambda_{2}(\mathbbm{1})p^{*}$.
The constraint $\lambda_{2}(w)\geq 1$ can be modeled by $L_{w}\succeq I-(1/n)J$ (the eigenvalues of $I-(1/n)J$ are $1,1,\ldots,1,0$) and (9) is the semidefinite program (SDP):
@@
SDPs are linear programs in the space of symmetric matrices where we are optimizing a linear function over a (convex) region described by linear matrix inequalities, or equivalently, an affine slice of the cone of psd matrices, see [4]. Calling (10) the primal SDP, its dual SDP is derived in [27] as
@@
As before, $w^{*}$ is optimal for (15) if and only if $\lambda_{n}(\mathbbm{1})w^{*}$ is optimal for (14) and $\lambda_{n}^{*}=\lambda_{n}(\mathbbm{1})q^{*}$. If $w^{*}$ is optimal for (15) then $\lambda_{n}(w^{*})=1$. Problem (15) is the SDP:
@@
We can also phrase this observation in terms of the dual SDPs:
@@
where and $u$ is a normalized eigenvector of $\lambda_{2}(\mathbbm{1})$ (respectively, $\lambda_{n}(\mathbbm{1})$). These observations were noted in [27]. A very short explanation why the notion of conformal rigidity simplifies when the eigenspace associated to $\lambda_{2}$ (or $\lambda_{n}$) has multiplicity 1 is as follows: in that case, we have
@@ 3.2. Symmetry reduction
(24) $\overline{w^{\ast}}:=\frac{1}{\textup{Aut}(G)}\sum_{\pi}\pi\cdot w^{\ast}\in\mathcal{F}.$
@@ 4. Certificates and Spectral Embeddings
Once can check that $\|q_{i}-q_{j}\|=2$ for all $ij\in E$, while if $ij\not\in E$, then $\|q_{i}-q_{j}\|=0,2$. For instance, $(1,9),(1,10)\in E$ and $(1,2),(1,16)\not\in E$. Therefore, there are two different distances between non-adjacent vertices of $G$ in the embedding provided by $\mathcal{Q}$. Since the eigenspace $\mathcal{E}_{8}$ is $1$-dimensional, any other embedding on this eigenspace must be a scaling of the configuration $\mathcal{Q}$. The scaling cannot change that there will be two different distances between non-adjacent vertices. ∎
@@
(27) $X_{ii}+X_{jj}-2X_{ij}=1\forall ij\in E,\,\,\,\textup{cols}(X)\in\mathcal{E}_{\lambda_{2}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}X\mathbbm{1}=0$
(28) $Y_{ii}+Y_{jj}-2Y_{ij}=1\forall ij\in E,\,\,\,\textup{cols}(Y)\in\mathcal{E}_{\lambda_{n}(\mathbbm{1})}\,\,\,\mathbbm{1}^{\top}Y\mathbbm{1}=0$
(29) $X^{\prime}_{ii}+X^{\prime}_{jj}-2X^{\prime}_{ij}=1\forall ij\not\in E,\,\,\,\textup{cols}(X^{\prime})\in\mathcal{E}_{\lambda_{n}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}X^{\prime}\mathbbm{1}=0$
(30) $Y^{\prime}_{ii}+Y^{\prime}_{jj}-2Y^{\prime}_{ij}=1\forall ij\not\in E,\,\,\,\textup{cols}(Y^{\prime})\in\mathcal{E}_{\lambda_{2}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}Y^{\prime}\mathbbm{1}=0.$
@@
A class of graphs that have edge-nonedge-isometric embeddings on $\mathcal{E}_{\lambda_{2}(\mathbbm{1})}$ and $\mathcal{E}_{\lambda_{n}(\mathbbm{1})}$ are the connected strongly-regular graphs. If $G$ is a connected strongly-regular graph then there are only two distances among pairs of vertices in $G$; $d(i,j)=1$ if $ij\in E$ and $d(i,j)=2$ if $ij\not\in E$. Also $G$ has only two nonzero eigenvalues which are typically called $r$ and $s$. Complements of strongly regular graphs are also strongly-regular and suppose $r^{\prime}=n-s$ and $s^{\prime}=n-r$ are the nonzero eigenvalues of $G^{c}$. We will see in Corollary 5.4 that $G$ has a spherical edge-isometric embedding $\mathcal{U}=\{u_{1},\ldots,u_{n}\}$ on $\mathcal{E}_{r}$ where $u_{i}^{\top}$ form the rows of a matrix $U$ whose columns form an orthonormal basis of $\mathcal{E}_{r}$. Since $\mathcal{E}_{r}=\mathcal{E}^{c}_{s^{\prime}}$, $\mathcal{U}$ is also a spherical embedding of $G^{c}$ on $\mathcal{E}^{c}_{s^{\prime}}$ and by Corollary 5.4, it is an edge-isometric embedding of $G^{c}$. Since the edges of $G^{c}$ are the nonedges of $G$ and vice-versa, $\mathcal{U}$ is an edge-nonedge-isometric embedding of $G$ on $\mathcal{E}_{r}$ (and of $G^{c}$ on $\mathcal{E}^{c}_{s^{\prime}}$). Similarly, there is an edge-nonedge-isometric spherical embedding $\mathcal{V}$ of $G$ on $\mathcal{E}_{s}=\mathcal{E}^{c}_{r^{\prime}}$ (and $G^{c}$ on $\mathcal{E}^{c}_{r^{\prime}}$). The embeddings $\mathcal{U}$ and $\mathcal{V}$ are spherical $2$-designs [11, Corollary 6.2]. Lemma 6.3 in [11] says that if $\mathcal{U}$ is a spherical $2$-design such that $u_{i}^{\top}u_{j}\in\{\alpha,\beta\}$ whenever $u_{i}\neq u_{j}$, and $G$ is the graph with vertex set $\mathcal{U}$ and two vectors adjacent if their inner product is $\alpha$, then $G$ is strongly- regular or complete.

### v2

@@ Abstract.
Conformal rigidity requires an extraordinary amount of structure in $G$. Every edge-transitive graph is conformally rigid. We prove that every distance-regular graph, and hence every strongly-regular graph, is conformally rigid. Certain special graph embeddings can be used to characterize conformal rigidity. Cayley graphs can be conformally rigid but need not be, we prove a sufficient criterion. We also find a small set of conformally rigid graphs that do not belong into any of the above categories; these include the Hoffman graph, the crossing number graph 6B and others. Conformal rigidity can be certified via semidefinite programming, we provide explicit examples.
@@ 1.2. Motivation
A final motivation comes from a recently proposed framework for Graph Sparsification due to Babecki and the authors [1]. In this framework, the hypercube graph cannot be further sparsified. We were wondering how general this result is which is what lead us to the notion of conformally rigid graphs. If (up to scaling) $\mathbbm{1}$ is the unique weight that maximizes $\lambda_{2}(w)$, then $G$ cannot be sparsified in any setting of [1]. This led us to wonder if all Cayley graphs have this property. As we will see in this paper, not all Cayley graphs have this property. This line of investigation led us to the notion of conformal rigidity. It is an interesting question to understand which conformally rigid graphs have unique maximizers for one or both of $\lambda_{2}(w)$ and $\lambda_{n}(w)$.
@@ 1.3.1. The optimization perspective
Remark. The spectral width minimization problem looks similar to conformal rigidity and so we pause to point out an important difference. In our setup we are solving
@@ 2.3. Distance-regular graphs
Note that, in particular, every strongly-regular graph is distance-regular. Strongly-regular graphs are distance-regular graphs of diameter two (where diameter is the largest distance between any pair of points). For example, the Shrikhande graph and its complement (see Fig. 1) are strongly-regular. The proof of Theorem 2.2 uses structural theory of distance-regular graphs (see [11]) and the duality theory of semidefinite programs which are developed in § 3 and § 4.
@@ 2.4. Cayley Graphs
A natural class of graphs with a great deal of symmetry are Cayley graphs. The Cayley graph $\Gamma(G,S)$ has vertex set $V=G$ where $G$ a finite group and the edge set $E=\left\{(g,s\circ g):g\in G,s\in S\right\}$ where $S\subset G$ is a generating set. Since we only deal with undirected graphs, we want the edge set to be symmetric which is achieved by requiring a symmetry in the generating set $S$, meaning $s\in S\implies s^{-1}\in S$. We always assume this.
Cayley graphs need not be conformally rigid. The smallest example we found is shown on the left in Fig. 5. This unweighted graph has $\lambda_{2}=2$ and $\lambda_{6}=5$. As it turns out, both of these eigenvalues can be improved by changing the edge weights: by increasing the weights on the three edges connecting the two triangles (while decreasing the weight on the other edges to maintain the normalization), we can increase $\lambda_{2}$ and attain $\lambda_{2}(w)=18/7>2$. Likewise, by setting the three edge weights connecting the two triangles to 0 (and increasing the remaining weights uniformly) we can decrease the largest eigenvalue to $\lambda_{6}(w)=4.5<5$ (and thus, in particular, can achieve values smaller than 5 for small positive weights by continuity of the eigenvalues). On the other hand, the circulant graph shown on the right of Fig. 5 is conformally rigid but not edge-transitive. Its conformal rigidity is implied by the following Theorem.
@@
Figure 5. Left: The triangular prism graph is Cayley and not conformally rigid. Right: a Cayley graph on $\mathbb{Z}_{18}$ (generated by $S=\left\{-5,-1,1,5\right\}$) that is conformally rigid and not edge-transitive.
@@
For the circulant graph on the right in Fig. 5, $V={\mathbb{Z}}_{18}=\{0,1,2,\ldots,17\}$ and $S=\{1,-1,5,-5\}$ we choose
@@ 2.5. Circulants
Denote a circulant with $n$ vertices and generators $S$ by $C_{n}(S)$. It is typical to list only one of $s$ or $-s$ in $S$, with the assumption that the other one is included. The set of vertices of $C_{n}(S)$ is $V={\mathbb{Z}}_{n}=\left\{0,1,2,\dots,n-1\right\}$ and its edges are $(v,v+s)$, as we vary over all $v\in V$ and $s\in S$ (mod $n$). Since circulants are Cayley graphs, Theorem 2.3 applies to them. We start with some examples.
@@
Circulants are rich and rather nontrivial. In particular, understanding when a circulant is edge-transitive appears to be a nontrivial problem, no definite characterization appears to be known. There are infinitely many circulants that are not conformally rigid.
@@
    Somewhat unrelated, but of independent interest in this context: is it possible to characterize when $C_{n}(\left\{a,b\right\})$ is edge-transitive in terms of $n,a,b$? Some partial results can be found in [2, 5, 6, 25, 28, 29].
@@ 2.6. Sporadic outliers
The graphs in this list are mostly well-known; the less well known examples (3), (5), (6), (8) are implemented in Mathematica 13.3 as (”Haar”, 565), (”ZeroTwoBipartite”, 7, 1),(”ZeroTwoBipartite”, 7, 2) and (”NoncayleyTransitive”, 24, 23). The graphs in Theorem 2.6 were all proven to be conformally rigid by identification of a suitable dual semidefinite programming certificate (see § 6). However, we also found one other example, the non-Cayley vertex-transitive graph (20,10) on $n=20$ vertices, that is not edge-transitive, not distance-regular, but very likely conformally rigid (beyond any reasonable numerical doubt); however, we were unable to find a certificate that would rigorously prove its conformal rigidity.
@@ 3. Semidefinite programming certificates for conformal rigidity
Assume that $G=([n],E)$ is an undirected connected graph with positive edge weights $w_{ij}$. Let $\mathcal{S}_{+}^{n}$ be the cone of $n\times n$ real symmetric psd matrices. We use $X\succeq 0$ to denote that $X\in\mathcal{S}^{n}_{+}$. For $A,B\in\mathcal{S}^{n}_{+}$, $A\succeq B$ stands for $A-B\succeq 0$.
@@
Problem (8) has an optimal solution since its objective function is continuous and its feasible region is compact. Therefore, (9) also has an optimal solution. If $w^{*}$ is optimal for (9), then $\lambda_{2}(w^{*})=1$ and hence there is no harm in optimizing over $\lambda_{2}(w)\geq 1$. The two problems are off by a scaling; $\lambda_{2}^{*}=|E|/p^{*}$ and $w^{*}$ is optimal for (9) if and only if $(|E|w^{*})/p^{*}$ is optimal for (8). Using $I$ to denote the identity matrix and $J$ to denote the matrix filled with 1’s, the constraint $\lambda_{2}(w)\geq 1$ can be modeled by $L_{w}\succeq I-(1/n)J$ (the eigenvalues of $I-(1/n)J$ are $1,1,\ldots,1,0$) and (9) is the semidefinite program (SDP):
@@
SDPs are optimization problems in the space of symmetric matrices where we are optimizing a linear function over a (convex) region described by linear matrix inequalities, or equivalently, an affine slice of the cone of psd matrices, see [4]. Calling (10) the primal SDP, its dual SDP is derived in [27] as
@@
As before, if $w^{*}$ is optimal for (15) then $\lambda_{n}(w^{*})=1$, $\lambda_{n}^{*}=|E|/q^{*}$ and $w^{*}$ is optimal for (15) if and only if $(|E|w^{*})/q^{*}$ is optimal for (14). Problem (15) is the SDP:
@@
Under strong duality, the pair of primal and dual SDPs will have the same optimal value, which allows the following re-phrasing:
@@
where $u$ is a normalized eigenvector of $\lambda_{2}(\mathbbm{1})$ (respectively, $\lambda_{n}(\mathbbm{1})$). These observations were noted in [27]. A very short explanation why the notion of conformal rigidity simplifies when the eigenspace associated to $\lambda_{2}$ (or $\lambda_{n}$) has multiplicity 1 is as follows: in that case, we have
@@ 3.2. Symmetry reduction
(24) $\overline{w^{\ast}}:=\frac{1}{|\textup{Aut}(G)|}\sum_{\pi}\pi\cdot w^{\ast}\in\mathcal{F}.$
@@ 4. Certificates and Spectral Embeddings
Once can check that $\|q_{i}-q_{j}\|=2$ for all $ij\in E$, while if $ij\not\in E$, then $\|q_{i}-q_{j}\|=0,2$. For instance, $(1,9),(1,10)\in E$ and $(1,2),(1,16)\not\in E$. Therefore, there are two different distances between non-adjacent vertices of $G$ in the embedding provided by $\mathcal{Q}$. Since the eigenspace $\mathcal{E}_{8}$ is $1$-dimensional, any other embedding on this eigenspace must be a scaling of the configuration $\mathcal{Q}$. The property of there being two different distances between non-adjacent vertices is invariant under scaling. ∎
@@
(27) $X_{ii}+X_{jj}-2X_{ij}=1\forall ij\in E,\,\,\,\textup{cols}(X)\subset\mathcal{E}_{\lambda_{2}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}X\mathbbm{1}=0$
(28) $Y_{ii}+Y_{jj}-2Y_{ij}=1\forall ij\in E,\,\,\,\textup{cols}(Y)\subset\mathcal{E}_{\lambda_{n}(\mathbbm{1})}\,\,\,\mathbbm{1}^{\top}Y\mathbbm{1}=0$
(29) $X^{\prime}_{ii}+X^{\prime}_{jj}-2X^{\prime}_{ij}=1\forall ij\not\in E,\,\,\,\textup{cols}(X^{\prime})\subset\mathcal{E}_{\lambda_{n}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}X^{\prime}\mathbbm{1}=0$
(30) $Y^{\prime}_{ii}+Y^{\prime}_{jj}-2Y^{\prime}_{ij}=1\forall ij\not\in E,\,\,\,\textup{cols}(Y^{\prime})\subset\mathcal{E}_{\lambda_{2}(\mathbbm{1})},\,\,\,\mathbbm{1}^{\top}Y^{\prime}\mathbbm{1}=0.$
@@
A class of graphs that have edge-nonedge-isometric embeddings on $\mathcal{E}_{\lambda_{2}(\mathbbm{1})}$ and $\mathcal{E}_{\lambda_{n}(\mathbbm{1})}$ are the connected strongly-regular graphs. If $G$ is a connected strongly-regular graph then there are only two distances among pairs of vertices in $G$; $d(i,j)=1$ if $ij\in E$ and $d(i,j)=2$ if $ij\not\in E$. Also $G$ has only two nonzero eigenvalues which are typically called $r$ and $s$. Complements of strongly regular graphs are also strongly-regular and suppose $r^{\prime}=n-s$ and $s^{\prime}=n-r$ are the nonzero eigenvalues of $G^{c}$. We will see in Corollary 5.4 that $G$ has a spherical edge-isometric embedding $\mathcal{U}=\{u_{1},\ldots,u_{n}\}$ on $\mathcal{E}_{r}$ where $u_{i}^{\top}$ form the rows of a matrix $U$ whose columns form an orthonormal basis of $\mathcal{E}_{r}$. Since $\mathcal{E}_{r}=\mathcal{E}^{c}_{s^{\prime}}$, $\mathcal{U}$ is also a spherical embedding of $G^{c}$ on $\mathcal{E}^{c}_{s^{\prime}}$ and by Corollary 5.4, it is an edge-isometric embedding of $G^{c}$. Since the edges of $G^{c}$ are the nonedges of $G$ and vice-versa, $\mathcal{U}$ is an edge-nonedge-isometric embedding of $G$ on $\mathcal{E}_{r}$ (and of $G^{c}$ on $\mathcal{E}^{c}_{s^{\prime}}$). Similarly, there is an edge-nonedge-isometric spherical embedding $\mathcal{V}$ of $G$ on $\mathcal{E}_{s}=\mathcal{E}^{c}_{r^{\prime}}$ (and $G^{c}$ on $\mathcal{E}^{c}_{r^{\prime}}$). The embeddings $\mathcal{U}$ and $\mathcal{V}$ are spherical $2$-designs [11, Corollary 6.2]. Lemma 6.3 in [11] says that if $\mathcal{U}$ is a spherical $2$-design such that $u_{i}^{\top}u_{j}\in\{\alpha,\beta\}$ whenever $u_{i}\neq u_{j}$, and $G$ is the graph with vertex set $\mathcal{U}$ and two vectors adjacent if their inner product is $\alpha$, then $G$ is strongly-regular.